OpenStudy (anonymous):

The solubility of acetanilide in hot water (5.5 g/100 ml at 100 C) is not very great, and it has some solubility in cold water (0.53 g/100 ml at 0 C). What would be the maximum theoretical percent recovery from the crystallization of 4.0 g of acetanilide from 100 ml of water (assuming the solution is chilled to 0 C for filtration). Show your calculations.

OpenStudy (paxpolaris):

\[\large percent\ change ={New-Old \over Old}\ \times100\%\] for calculating the percent recovery... ignore the negative sign of the above formula. New= the final mass of acetanilide in 100ml water at 0°C = 0.53g Old= mass in original 100ml solution = 4.0g \[{\left| 0.53-4.0 \right| \over 4.0} \times 100\% \\={3.47\over4.0}\times 100\% \\ \Large \color{green}{=87\%}\]